OpenStudy (anonymous):

implicit differentiation y^2 +4xy +x^2 = 13... i need to know how the 4xy gets differentiated with product rule

OpenStudy (anonymous):

4y+4x(dy/dx)

OpenStudy (anonymous):

not correct.. i am pretty sure you are wrong

OpenStudy (anonymous):

think of \(y=f(x)\) so every place you see a \(y\), mentally replace it by \(f(x)\)

OpenStudy (anonymous):

i have (4) (xy) so.... 4' times xy + 4 times xy'

OpenStudy (anonymous):

\[y^2 +4xy +x^2 = 13..\] \[f^2(x)+4xf(x)+x^2=13\] \[2f(x)f'(x)+4(f(x)+xf'(x))+2x=0\] more easiliy written as \[2yy'=4(y+xy')+2x=0\]

OpenStudy (amistre64):

lol, if anything [4xy]' = 4'xy + 4x'y + 4xy'

OpenStudy (anonymous):

the 4 is a constant, so treat it like one you don't use the product rule when you take the derivative of \(4\sin(x)\) you just write \(4\cos(x)\) right?

OpenStudy (amistre64):

you can do the product rule on it; the 4' part goes 0 tho

OpenStudy (anonymous):

the product rule is needed because you have \(x\times y\) not because of the 4

OpenStudy (anonymous):

i dont know .. im still confused.. im having a moment.. ive done these before.. but i dont understand right now

OpenStudy (anonymous):

ok suppose you had this: \[x\sin(x)\] what would you do to find the derivative?

OpenStudy (anonymous):

you have to use the product rule, because it is a product. you have to write \[\sin(x)+x\cos(x)\]

OpenStudy (anonymous):

now suppose instead you had \[4x\sin(x)\] the 4 doesn't bother you at all, it is still a constant, so the derivative would be \[4(\sin(x)+x\cos(x))\] the product rule is still necessary, but not because of the 4, but because you have x times sine

OpenStudy (anonymous):

now instead of \(x\sin(x)\) you have \(xy\) where \(y\) is some unknown function of \(x\) a

OpenStudy (anonymous):

product rule still applies

OpenStudy (anonymous):

but the 4 is just sitting there, so the derivative of \(4xy\) with respect to \(x\) is \[4(y+xy')\]

OpenStudy (anonymous):

so 4xy' is 4y +4xy'

OpenStudy (anonymous):

yes that is the deriviative, right

OpenStudy (anonymous):

hope that is more or less clear

OpenStudy (anonymous):

its better now.. thanks

OpenStudy (anonymous):

yw